# Consecutive Model Editing with Batch alongside HooK Layers

### v1

@@ 4.1 Experiment Setups
We employ three popular editing evaluation metrics defined in (Yao et al., 2023; Huang et al., 2023; Cao et al., 2021), i.e., reliability, generality, and locality. Given an initial base model $f_{\theta}$, a post-edit model $f_{\theta^{\prime}}$, and a set of edit instances $\{(x_{t},y_{t})\}$, the reliability is computed as the average accuracy of the edit cases:
@@
Table 2: Consecutive batch editing results with batch size 10 on 1k samples.
Method Model ZsRE COUNTERFACT
Reliability Generality Locality Reliability Generality Locality
FT (Meng et al., 2022) GPT2-XL 0.0379 0.0248 0.0660 0.0100 0.0100 0.0600
MEND (Mitchell et al., 2022a) 0.0002 0.0001 0.0012 0.0001 0.0000 0.0008
MEMIT (Meng et al., 2023) 0.3488 0.3296 0.7074 0.5600 0.3700 0.3100
COMEBA-HK 0.6691 0.5611 0.9723 0.8600 0.3800 0.5900
FT (Meng et al., 2022) GPT-J 0.2353 0.2170 0.5527 0.0200 0.0200 0.7200
MEND (Mitchell et al., 2022a) 0.0005 0.0001 0.0002 0.0100 0.0100 0.0100
MEMIT (Meng et al., 2023) 0.6336 0.4890 0.7480 0.7500 0.4500 0.4200
COMEBA-HK 0.7989 0.6129 0.9652 0.9500 0.4100 0.8000
4.2 Evaluation on Single-round Batch Editing
We first test the effectiveness of our method under basic single-round batch editing settings, i.e., the model is rolled back to the initial state after each batch editing. Both MEMIT and COMEBA-HK need to set the parameter $\lambda$, the balance factor between pre-trained and newly updated associations. According to (Meng et al., 2023), higher $\lambda$ helps preserve the original model behavior (locality) but could be harmful to reliability and generality, and the best overall performance is found at around $\lambda=10^{4}$. However, with the intent to verify whether our method comprehensively improves the editing, that is, could accept lower $\lambda$ to assign higher weight for new associations while not sacrificing the locality, we deliberately set $\lambda=5\times 10^{3}$ for COMEBA-HK and keep it as the optimized value for MEMIT, which are $2\times 10^{4}$ and $1.5\times 10^{4}$ for GPT2-XL and GPT-J respectively.
The evaluation results are shown in Table 1. COMEBA-HK outperforms all baselines except the reliability metric for GPT-J. Specifically, despite the relatively low $\lambda$, our method overwhelms other baselines in reliability and generality metrics for GPT2-XL meanwhile maintaining a better locality. This indicates that lowering $\lambda$ or, in other words, increasing the weight of the new associations does not sacrifice the locality in COMEBA-HK, which demonstrates the all-sided improvement of our method. The improvement margin is less for GPT-J compared with those for GPT2-XL, yet our method still ranks first for most of the metrics in both datasets. Interestingly, the reliability in the COUNTERFACT dataset for GPT-J is lower than that of MEMIT and even lower than the performance of GPT2-XL. We believe this may be one side-effect of descending $\lambda$, which inevitably gives rise to the increase of weight change magnitude at each batch editing step and consequently distorts previously updated associations more severely (sec.4.5). However, this phenomenon does not apply to consecutive editing cases, which is shown in the next section.
4.3 Evaluation on Consecutive Batch Editing
We evaluate our method’s capability on 1k samples from both datasets for consecutive batch editing, i.e., there is no roll-back. The evaluation is conducted after the end of the whole consecutive batch editing process. We set $\lambda$ to $15,000$ as the scenario now is consecutive batch editing and set the batch size to 10, that is, 100 batch editing steps.
Results are shown in Table.2. Our method surpasses other methods in 100 consecutive steps with an even larger improvement margin for almost all the metrics compared to the single-round batch editing. This demonstrates that our method does not depend on simple regurgitation of the editing samples nor rely heavily on the trade-offs of lowering the balancing factor $\lambda$ to increase the reliability and locality.
We extend the consecutive steps to 1k to explore the limit of our method. Results can be found in Fig.7. Surprisingly, the locality experiences a great fall from 1k to 2k steps but remains steady from 200 to 1k editing steps, which proves that the hook layer stably obstructs the out-scope samples. Reliability and generality consistently fall as the consecutive steps grow, indicating that there is still room for improvement in this field.
4.4 Validation of Local Editing Scope
Figure 3: Difference in z-score between updated keys and others.
Given an updated hook layer with the weight $W_{h}$, the original model weight $W_{0}$, an updated key $k_{i}$, and an out-of-scope key $k_{j}$, we conduct experiments to verify whether $\parallel{W_{h}k_{i}-W_{0}k_{i}}\parallel\gg\parallel{W_{h}k_{j}-W_{0}k_{j}}\parallel$ holds. We select 100 samples from the COUNTERFACT dataset to edit GPT2-XL using COMEBA-HK, then apply the edited model to these 100 samples and record the z-score to the L2-norm of the difference vector between the last hook layer response and original model response of the updated key (the last subject token), namely, z-score of $\parallel{W_{h}k_{i}-W_{0}k_{i}}\parallel$. Both reliability and generality prompts are included for comprehensiveness.
The result is shown in Fig.3. Almost all the z-scores of the responses from updated keys exhibit a great margin from the mean value, with the lowest around 1.5 in reliability samples and 2 in generality samples. This not only validates our editing scope identification (section 3.1.3) technique but also cleans a possible doubt about the undesired block of generality instances in the hook layer. In addition, the discriminative z-score could also ensure the preciseness of the swapping, which can improve the locality.
4.5 Detailed Analysis and Discussions
We further conduct ablation studies and detailed analyses of different settings of the proposed method. We investigate cases with different sample sizes that correspond to different numbers of consecutive steps for each experiment. Note that the evaluation is conducted after the corresponding entire consecutive steps are finished. For example, in Fig.4, we conduct experiments for sample sizes 200, 400, 600, 800, and 1000, so the evaluation is triggered right after the first 200, 400, 600, etc, samples are edited to the model. Unless specified, the batch size is selected to be 10.
Ablation Study of Update Mechanism and Hook Layers
Figure 4: Ablation study for update mechanism and hook layers.
The effectiveness of the derived consecutive updating mechanism and the hook layers are discussed in this part. We run three cases using GPT2-XL, namely, MEMIT (no consecutive updating mechanism, no hook layers), COMEBA without hook (COMEBA w/o HK), and COMEBA-HK for consecutive batch editing on 1k samples from both ZsRE and COUNTERFACT datasets.
The results are demonstrated in Fig.4. In almost all metrics of the two datasets except the generality of COUNTERFACT, the COMEBA w/o the hook performs better than the vanilla MEMIT, and the margin tends to increase as the consecutive steps ascend. This certifies the effectiveness of our derived consecutive updating mechanism in consecutive batch editing scenarios. For the ZsRE dataset, the method with hook layers considerably outperforms the one without hook in the locality without sacrificing reliability and generality. This verifies that the hook layer can efficiently and accurately block the out-scope instances from the input without fraudulently missing in-scope instances. For the COUNTERFACT dataset, the reliability of COMEBA-HK is consistently higher than the other two, and the generality surpasses that of MEMIT after 80 editing steps. It seems that the hook layer causes some side effects in the locality of COUNTERFACT, but this circumstance is not found in the ZsRE dataset. It is worth noting that COMEBA-HK shows the most stable performance as the number of consecutive editing steps grows, demonstrating the great potential of our method for consecutive editing.
Effect of the Number of Editing Layers
Figure 5: Performance comparisons on different number of editing layers. Layers are selected from the critical path identified in (Meng et al., 2023).
To investigate the necessity of applying the hook layer onto multiple transformer layers, we conduct the consecutive batch editing experiment on the ZsRE dataset for GPT2-XL (Fig.5). As the effect of choosing different layers has already been studied in (Meng et al., 2023), we focus only on the effect of the number of layers. We selected the last one, three, and all number of layers from the critical path identified in (Meng et al., 2023; Yao et al., 2023), respectively.
Effect of the Balance Factor $\lambda$
Figure 6: Performance comparisons on different values of $\lambda$ on 1k samples.
Figure 7: Performance comparisons on different values of $\lambda$ on 10k samples.
We test the effect of different $\lambda$, the balance factor between pre-training and newly updated associations. We first evaluate the COMEBA-HK with different $\lambda$ on 1k samples from ZsRE (Fig.6). It seems that a small value of $\lambda=1,000$ would cause significant damage to all three metrics, especially the reliability and generality, since they experience a great drop as the consecutive steps increase. This may result from the overly high magnitude of the weight change caused by the low value of $\lambda$, which severely distorts the previously updated associations. Meanwhile, a too-high value of $\lambda=20,000$ seems also not to be a good choice as well, which gives rise to an overly small magnitude of the weight change so that it fails to deliver the new optimized values for keys. The cases of $\lambda=5000,10000,15000$ do not show an apparent difference, so we extend further the sample size to 10k (Fig.7).
Extended results show that 5000 is not a good choice for large-consecutive editing steps, though it performs no worse than the other two in early 1k samples. The case of $\lambda=15,000$ ranks first in reliability and generality. Although it has worse performance in locality compared to $\lambda=10,000$, the margin between them gradually narrows as the consecutive steps rise. Overall, we believe 15,000 would be a reasonable selection.
5 Related Work
Recent years have witnessed prosperous development in the field of model editing. According to (Yao et al., 2023), the proposed methods so far can be generally classified into two groups, i.e., those directly modify the model’s weight, and others that do not.
The methods that do not directly alter the model weights generally follow two directions: they either employ an external memory or introduce additional adjustable parameters. Methods like T-Patcher (Huang et al., 2023) and CaliNET (Dong et al., 2022) apply new neurons that are responsible for specific mistakes in the last layer of the FFN model. Grace (Hartvigsen et al., 2022) introduces a timely adjusted code book to edit the model’s behavior. Another group of methods like (Mitchell et al., 2022b) integrates an explicit external memory as edit descriptors to help editing scope recognition.
Those directly altering the model’s weight either train a hyper-network to predict the change required by the edits (Mitchell et al., 2022a; De Cao et al., 2021) or first locate corresponding parameters that are responsible for specific knowledge and then edit the located parameters (Meng et al., 2023, 2022; Li et al., 2023; Dai et al., 2022).
6 Conclusion
In this work, we introduce a novel model editing method, COMEBA-HK, which advocates the more practical consecutive batch model editing. COMEBA-HK uses an expanded editing mechanism to support consecutive editing and newly proposed hook layers to identify the editing scope. Compared to existing model editing methods, COMEBA-HK does not require large external memory nor extra training for meta-networks or classifiers. It only needs an extra tiny amount of memory to store the edited hook layer. Comprehensive experiments are conducted to verify the effectiveness of the method over both single-round and consecutive batch editing. We also investigate the validity of the editing scope recognition in hook layers and the influence of each parameter in our method.
Impact Statements
The goal of our method is to conveniently add or fix knowledge in the LLMs without expensive retraining so that time and resources can be saved. However, we caution readers that our method should not be used to edit any malicious or offensive knowledge, including but not limited to, political authoritative facts, rumors, ethical discrimination statements, etc. It is worth reminding that the safe and responsible application of our method is really important. None of offensive, toxic, or malicious editing should be allowed.
Acknowledgements
Do not include acknowledgements in the initial version of the paper submitted for blind review.
If a paper is accepted, the final camera-ready version can (and probably should) include acknowledgements. In this case, please place such acknowledgements in an unnumbered section at the end of the paper. Typically, this will include thanks to reviewers who gave useful comments, to colleagues who contributed to the ideas, and to funding agencies and corporate sponsors that provided financial support.
Appendix
@@ Acknowledgements
Appendix A Dataset and Implementation Details
Datasets
@@ Appendix A Dataset and Implementation Details
We generally follow the procedure conducted in (Meng et al., 2023, 2022). We fine-tune the $mlp_{proj}$ parameter from layer 0 for GPT2-XL and layer 21 for GPT-J since they are found to have the optimal performance. For both models, we conduct 25 optimization steps using Adam optimizer (Kingma & Ba, 2015) and use learning rate $5e^{-4}$. All other parameters of both models follow default settings.
@@
MEND (Mitchell et al., 2022a) conducts the editing by manipulating the language models’ gradient. It trains a meta-network that employs a rank-1 decomposition of the model gradients and predicts a new rank-1 update to the corresponding model weights. In this work, we train two meta-networks using corresponding training split in the ZsRE and COUNTERFACT datasets for GPT2-XL following the default hyperparameter settings. Due to the large required computation resource for training GPT-J (6B) meta-network, we do not perform training for GPT-J. Instead, we directly download the available MEND weights for GPT-J provided by (Meng et al., 2022) and use that for the evaluation of MEND.
@@
Mitchell et al. (2022b) designed a memory-augmented editing method, which requires an external cache to store explicit editing cases. It also adopts a scope classifier that determines whether an input sample falls in the editing scope and a small counterfactual model for editing the in-scope cases. For both GPT2-XL and GPT-J, we train two separate models for the two datasets respectively. Following the original paper, we apply distilbert-base-cased (Sanh et al., 2019) for the scope classifier across all models. For the small counterfactual model, we employ GPT2 for GPT2-XL and gpt-j-tiny-random [^2] https://huggingface.co/anton-l/gpt-j-tiny-random for GPT-J. All hyperparameters follow default settings.
@@
MEMIT (Meng et al., 2023) treats the feedforward layer of transform as a linear associative memory and uses a minimum square error optimization to add new key-value associations to layer weights. We follow the original paper to edit the layers that are in the identified critical path and set the balance factor $\lambda$ to the optimal value found in the original work. Other parameters for the two models are all set based on configurations in (Meng et al., 2023, 2022).
COMEBA-HK
COMEBA-HK expands the update mechanism in MEMIT to consecutive cases and applies hook layers to separate the weight change from the original model layer. For both models, we set $\lambda=15,000$, $\alpha_{z}=2.2$. Unless specified, we evaluate our method on full critical path layers identified in (Meng et al., 2023). We employ the same procedure in MEMIT (Meng et al., 2023) to compute the updating keys and the target values, except that the most recently updated model during the process of consecutive editing is applied for relevant computations. We applied ”torch.float16” for the GPT-J model for all experiments.
Appendix B Analysis of Hyperparameters
@@
To verify the significance of the dynamical determination process, we also test the fix $\alpha$ case. We chose the value of 3 as it is the standard threshold used in standardization to detect outliers. The results reveal a dramatic decline in reliability and generality and perfect fulfillment in the locality, indicating the fact that almost all of the instances are indiscriminately obstructed by the hook layers regardless of the editing scope. Besides, it is practically unrealistic to choose an optimal fixed $\alpha$ before the editing process. Therefore, we believe it is more reasonable to decide $\alpha$ dynamically.
Effect of Editing Batch Size
Figure 11: Performance comparisons on different editing batch sizes on 1k samples of ZsRE.
The results showcase that while a decrease in the three metrics exists as the batch size rises, the margin could be negligible, denoting that our method possesses the mass-editing capacity.

### v3

@@ Appendix A Experiment Details
We employ three popular editing evaluation metrics defined in Yao et al. (2023); Huang et al. (2023); Cao et al. (2021), i.e., reliability, generality, and locality. Given an initial base model $f_{\theta}$, a post-edit model $f_{\theta^{\prime}}$, and a set of edit instances $\{(x_{t},y_{t})\}$, the reliability is computed as the average accuracy of the edit cases:
@@
Datasets
@@
ZsRE is a question-answering dataset that uses back-translation to generate equivalent neighborhoods. It is initially used in factual knowledge evaluation and later adopted in model editing by Mitchell et al. (2022a). COUNTERFACT is a challenging dataset focusing on counterfactual information with a low prediction score compared to correct facts. It builds out-of-scope data by replacing the subject entity with a similar description that shares the same predicate.
@@ Appendix B Baselines and Implementation Details
We implemented two fine-tuning methods in the experiments. For FT-L, we follow the procedure in Meng et al. (2023, 2022) and fine-tune the $mlp_{proj}$ parameter from layer 0 for GPT2-XL and layer 21 for GPT-J since they are found to have the optimal performance. FT-M [^5] https://github.com/zjunlp/EasyEdit/blob/main/hparams/FT/gpt2-xl.yaml is a small variation of FT-L, and it uses a different loss computation procedure to optimize the parameters. For both models, we conduct 25 optimization steps using Adam optimizer Kingma and Ba (2015) and use learning rate $5e^{-4}$. All other parameters of both models follow default settings.
LoRA
Hu et al. (2022) proposed a parameter-efficient fine-tuning method that decomposes the update gradient matrix into two small rank-n matrices, which reduces the required memory for LLM training to a large extent. In all experiments, we set the learning rate and the rank number to $1e^{-3}$ and 8, respectively. The $\alpha$ was chosen to be 32, and the dropout rate was 0.1. The number of update steps is 30 for GPT2-XL and 50 for GPT-J.
@@
MEND Mitchell et al. (2022a) conducts the editing by manipulating the language models’ gradient. It trains a meta-network that employs a rank-1 decomposition of the model gradients and predicts a new rank-1 update to the corresponding model weights. In this work, we train two meta-networks using corresponding training split in the ZsRE and COUNTERFACT datasets for GPT2-XL following the default hyperparameter settings. Due to the large required computation resource for training GPT-J (6B) meta-network, we do not perform training for GPT-J.
@@
Mitchell et al. (2022b) designed a memory-augmented editing method, which requires an external cache to store explicit editing cases. It also adopts a scope classifier that determines whether an input sample falls in the editing scope and a small counterfactual model for editing the in-scope cases. For both GPT2-XL and GPT-J, we train two separate models for the two datasets, respectively. Following the original paper, we apply distilbert-base-cased Sanh et al. (2019) for the scope classifier across all models. For the small counterfactual model, we employ GPT2 for GPT2-XL and gpt-j-tiny-random [^6] https://huggingface.co/anton-l/gpt-j-tiny-random for GPT-J. All hyperparameters follow default settings.
@@
MEMIT Meng et al. (2023) treats the feedforward layer of transform as a linear associative memory and uses a minimum square error optimization to add new key-value associations to layer weights. We follow the original paper to edit the layers in the identified critical path and set the balance factor $\lambda$ to the optimal value found in the original work. Other parameters for the two models are all set based on configurations in Meng et al. (2023, 2022).
GRACE
Hartvigsen et al. (2022) proposed an editing method that preserves the original model parameters and adopts a codebook maintained by adding, splitting, and expanding keys over time to store relevant edits. We follow the optimized settings in the original paper and set the value optimizing learning rate to 1. The number of iterations for optimizing the values is 100, and the initial $\varepsilon$ value is chosen to be 1. The codebook is employed at layers 35 and 25, respectively.
CoachHooK
CoachHooK expands the update mechanism in MEMIT to consecutive cases and applies hook layers to separate the weight change from the original model layer. For both models, we set $\lambda=15,000$, $\alpha_{z}=2.2$ for consecutive batch editing. Unless specified, we evaluate our method on full critical path layers identified in Meng et al. (2023). We employ the same procedure in MEMIT Meng et al. (2023) to compute the updating keys and the target values, except that the most recently updated model during the process of consecutive editing is applied for relevant computations. We applied "torch.float16" for the GPT-J model for all experiments.
Appendix C More detailed analysis and discussions
Effect of the Number of Editing Layers
Figure 10: Performance comparisons on the different number of editing layers. Layers are selected from the critical path identified in Meng et al. (2023).
To investigate the necessity of applying the hook layer onto multiple transformer layers, we conduct the consecutive batch editing experiment on the ZsRE dataset for GPT2-XL (Fig. 10). As the effect of choosing different layers has already been studied in Meng et al. (2023), we focus only on the effect of the number of layers. We selected the last one, three, and all layers from the critical path identified in Meng et al. (2023); Yao et al. (2023), respectively.
@@ Appendix C More detailed analysis and discussions
To verify the significance of the dynamical determination process, we also test the fix $\alpha$ case. We chose the value of 3, the standard threshold used in standardization to detect outliers. The results reveal a dramatic decline in reliability and generality and perfect fulfillment in the locality, indicating that almost all instances are indiscriminately obstructed by the hook layers regardless of the editing scope. Besides, choosing an optimal fixed $\alpha$ before editing is practically unrealistic. Therefore, it would be more reasonable to decide $\alpha$ dynamically.
Model Type Inference Time (s) GPT2-XL Pre-edit 0.1187 Post-edit 0.1297 GPT-J Pre-edit 0.0762 Post-edit 0.0863
Table 3: Inference time analysis.
Model Granularity Instances Reliability Generality Locality GPT-J Instances 99.00 97.50 - Overall tokens 9.69 12.51 11.19 Unwanted tokens 0.38 0.13 11.19
Table 4: Percentage of instances/tokens that used the hook layer.
Deferral Radius Model COUNTERFACT Reliability Generality Locality Average $\varepsilon=1$ GPT2-XL 100 0.40 100 66.80 $\varepsilon=3$ 100 0.42 100 66.81 $\varepsilon=5$ 100 0.65 99.50 66.72 $\varepsilon=10$ 100 1.80 93.70 65.17 $\varepsilon=20$ 100 18.30 56.60 58.30 $\varepsilon=30$ 100 83.90 7.40 63.77 $\varepsilon=1$ GPT-J 100 0.50 100 66.83 $\varepsilon=3$ 100 0.54 100 66.85 $\varepsilon=5$ 100 0.57 100 66.86 $\varepsilon=10$ 100 0.68 99.60 66.76 $\varepsilon=20$ 100 5.00 93.20 66.07 $\varepsilon=30$ 100 31.30 58.90 63.40
Table 5: Results of GRACE with increased $\varepsilon$.
Investigation on hook layer employment
Although the validation of the hook layer has been proved in section 4.4, we conducted extra experiments to survey how many entries that should apply the hook layer indeed use the hook layer and vice versa. We investigated three granularity: instances, overall tokens, and unwanted tokens (Table 4). Suppose the number of instances is $A$, the total number of tokens for the set of instances is $T$, and there are $T^{\prime}$ tokens that used the hook layer and $A^{\prime}$ instances have their updated keys [^7] Each instance only has one updated key. (the last subject token) use the hook layer. The instance granularity was measured by $\frac{A^{\prime}}{A}$, the overall tokens granularity was calculated by $\frac{T^{\prime}}{T}$, and the unwanted tokens $\frac{T^{\prime}-A^{\prime}}{T}$.
The results show that almost all reliability and generality instances apply the hook layer, and few unwanted tokens mistakenly use the hook layer. This again demonstrates the effectiveness of our method’s editing scope identification.
GRACE with greater deferral radius
Although we followed the settings found in the original paper of GRACE Hartvigsen et al. (2022), one may argue that the terrible generality performance of GRACE in Table 2 is caused by the over small deferral radius ($\varepsilon$) and increasing it may help the model reach a better balance between generality and locality, then resulting in an improved overall average. Therefore, we further conducted the consecutive batch editing experiments for GRACE with several increased $\varepsilon$ on the COUNTERFACT dataset, the result is shown in Table 5.
It is not hard to find from the results that, though the results indeed show the trade-off between generality and locality, the average does not show great improvement. This proves that merely increasing the deferral radius for GRACE does not necessarily improve its overall average performance.
Inference Time Analysis
As our method will introduce new hook layers to the model, we conduct an experiment to investigate its influence on the model inference. We run GPT2-XL on NVIDIA Titan GPU and GPT-J on NVIDIA A6000. Table 3 shows the running result for the corresponding pre-edit and post-edit models. The hook layers’ employment does not seem to delay the model inference too much. This may result from the fact that the hook layers are only introduced for the small proportion of layers in the critical path, and the computation implemented in the hook layers is relatively simple.
Memory Analysis
Unlike GRACE Hartvigsen et al. (2022), whose memory requirement grows over time and SERAC Mitchell et al. (2022b), which needs extra memory for counterfactual model and scope classifier, the memory requirement of our method remains unchanged over time. Therefore, the final memory requirement is fixed no matter how many edits you make to the model. The initial memory requirement is acceptable since it is at maximum the copy of the 6 to 7 FFN projection layer weights in the model. Specifically, the hook layers are only applied to a set of identified layers, which usually accounts for a small proportion of the whole layers. For example, the number of identified layers for GPT-J-6B is 6, which is [3, 4, 5, 6, 7, 8], and 5 for GPT2-XL, which is [13, 14, 15, 16, 17]. Furthermore, it is not compulsory to hang hook layers to all the identified layers, user can decide how many layers they want to edit. For convenience, we assume to use all the identified layers here. Take the GPTJ-6B as an example, a projection FFN layer weight dimension is $16384\times 4096$, assuming the data type is float32, then GPU memory required by its hook layer (just a copy of itself) is approximately $\frac{16384\times 4096\times 4}{1024^{3}}=0.25$GB (ignore the bias). Now, the editing layers used in our approach for GPT-J-6B is [3,4,5,6,7,8], so the required memory is around $0.25\text{GB}\times 6=1.5\text{GB}$.